OpenStudy (anonymous):

I need help on the intergral of cos (pi/20) :S

OpenStudy (anonymous):

in your integration there is no variable please write the question properly.

OpenStudy (anonymous):

the full question is, pi/10 x cos (pi/20)

OpenStudy (anonymous):

i think \[\int\limits_{}^{}\cos \pi x/20dx\] let \[y=\pi x/20\] after this substitution you will get dx=20 dy/pi\[20/\pi \int\limits_{}^{}\cos y dy=-20\sin y/\pi =-(20/\pi)\sin (pix/20)\]

OpenStudy (anonymous):

sin(pi/20)

OpenStudy (anonymous):

see if there is no x under cosine function then it is simply a constant and you can take it outside the integration

OpenStudy (anonymous):

Thankyou so Much :D

OpenStudy (anonymous):

have you got your answer?

OpenStudy (anonymous):

Just doing it now (:

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

if you have any doubt please ask.

OpenStudy (anonymous):

Maths is hard.

OpenStudy (anonymous):

are you talking in general or you haven't understand any step of mine?

OpenStudy (anonymous):

General. Your steps helped alot.